# Composite Learning Backstepping Control With Guaranteed Exponential Stability and Robustness

### v3

@@ V-C Robustness Results
Theorem 3: For the system (4) with the additive disturbance $\bm{d}$ and Assumptions 1–3 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist suitable control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) has:
@@
    Partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$ when partial IE exists for constants $T_{\rm a}$, $\sigma\in\mathbb{R}^{+}$ and the set $\mathcal{I}$ no longer changes, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$;
@@
    Practical exponential stability on $t$ $\in$ $T_{\rm e},\infty)$ if IE exists for constants $T_{\rm e}$, $\sigma$ $\in$ $\mathbb{R}^{+}$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$.
@@ VI-A Stability and Convergence Comparisons
Set the control parameters $k_{{\rm c}i}=1$ with $i=1,2,3$, $\Gamma$ == $3I$, $\kappa=1$, $\tau_{\rm d}=3$, $\hat{\bm{\theta}}(0)=\bm{0}$, and $\sigma=10^{-4}$, and choose the stable filter $H(s)=25/(s^{2}+10s+25)$ in (16). Gaussian white noise with 0 mean and 0.001 standard deviation is added to the measurement of the states $x_{i}$. The MRE-HOT in [43] and the composite learning dynamic surface control (CL-DSC) in [49] are selected as baseline controllers, where we set the damping parameters $k_{\rm{d}\it i}=$ 0.1 for the MRE-HOT, the stable filter $L(s)$ == $20/(s+20)$ for the CL-DSC, and the other shared parameters to be the same values for fair comparisons.
@@
Case 1: Tracking with PE. Consider a tracking problem under PE with the reference trajectory $y_{\rm r}=1.5\sin(0.5t)$, the desired parameter $\bm{\theta}=[0.1,0.5,1.5]^{T}$, and the initial state $\bm{x}(0)$ == $0.6,0,0]^{T}$. Performance comparisons of the three controllers are depicted in Figs. 3. It is observed that the estimation error $\tilde{\bm{\theta}}$ by the CL-DSC has a steady-state error since the time derivative of each virtual control input is estimated by the filter $L(s)$ [see Fig. 3(a)], $\tilde{\bm{\theta}}$ by the MRE-HOT converges to $\bm{0}$ after running 30 s [see Fig. 3(a)] as PE is fulfilled in this case, and $\tilde{\bm{\theta}}$ by the proposed CLBC exhibits the rapid convergence to $\bm{0}$ [see Fig. 3(a)], which validates its strong learning capability. Besides, the proposed CLBC exhibits much better tracking performance than the CL-DSC and MRE-HOT [see Fig. 3(b)] because the exciting strength $\sigma_{\rm c}$ of the CLBC is monotonic non-decreasing and keeps a high level throughout [see Fig. 3(d)]. Moreover, the control inputs $u$ by the proposed CLBC and the MRE-HOT are not affected by the measurement noise throughout and are comparable after their estimation errors $\tilde{\bm{\theta}}$ converge to $\bm{0}$ [see Fig. 3(c)], but $u$ by the CL-DSC is seriously polluted by the measurement noise [see Fig. 3(c)], which results in much worse tracking performance [see Fig. 3(b)].
@@ VI-B Transient Performance Comparisons
with $u\in\mathbb{R}$, $\varphi_{1}(x_{1})=x_{1}^{2}$, $\theta=2$, and $\bm{x}(0)=[0.6,0]^{T}$. Note that for existing modular backstepping control methods, if the damping coefficients $k_{\rm{d}\it i}$ are close to 0, the transient performance of the above system can be deteriorated by the modeling error term $\Phi^{T}\tilde{\theta}$ regardless of the convergence of the estimation error $\tilde{\theta}$ [84]. The reference trajectories $y_{\rm r}$, $\dot{y}_{\rm r}$, and $\ddot{y}_{\rm r}$ are generated by $y_{\rm r}(t)$ == $\frac{a_{0}}{a(s)}[r(t)]$ with $a_{0}=1$, $a(s)=s^{3}+3s^{2}+3s+1$, and $r(t)=\sin(2t)$. Set the control parameters $k_{{\rm c}1}=k_{{\rm c}2}=1$, $\Gamma=1$, $\kappa=1$, $\tau_{\rm d}=1$, $\hat{\theta}(0)=0$, and $\sigma=10^{-4}$, and choose the stable filter $H(s)=25/(s+25)$. The HOT adaptive backstepping control (HOT-ABC) in [30] and the MRE-HOT in [43] are selected as baseline controllers, where their shared parameters are set to be the same for fair comparisons.
@@
Performance comparisons of transient tracking and parameter convergence of the three controllers are exhibited in Fig. 5, where both the damping parameters $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ increase from $0.01$ to $0.19$ with a step size of $0.03$, and the arrows indicate the increasing direction $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$. It is observed that the transient results of the HOT-ABC and MRE-HOT are deteriorated by decreasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ [see green and red dash lines in Fig. 5(a)]. For the MRE-HOT, increasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ reduces the initial oscillations of the tracking error $|e_{1}|$ [see green dash lines in Fig. 5(a)] but also slows down parameter convergence [see green dash lines in Fig. 5(b)]. The proposed CLBC provides the convergence of $|e_{1}|$ to 0 with the smallest initial oscillations [see Fig. 5(a)] and achieves the fastest parameter convergence [see Fig. 5(b)], which implies that: 1) the transient tracking and parameter convergence of the HOT-ABC and MRE-HOT are sensitive to damping parameters; 2) the transient performance of the proposed CLBC is enhanced by introducing the extra prediction error $\bm{\epsilon}$ in (22) without resorting to nonlinear damping terms; 3) The combination of two prediction errors $\bm{\xi}$ in (19) and $\bm{\epsilon}$ in (22) accelerates parameter convergence.
@@ VII Conclusions
This paper has presented a feasible modular backstepping control strategy named CLBC for strict-feedback uncertain nonlinear systems. The proposed composite learning HOT allows the exact implementation of the high-order time derivatives of parameter estimates and the offset of modeling errors, such that the transient performance can be guaranteed without resorting to nonlinear damping terms or high control gains. The proposed algorithm of staged exciting strength maximization ensures that the exciting strength is monotonously non-decreasing in each excitation stage, such that the exponential stability of the closed-loop system with parameter convergence is obtainable under the much weaker condition of IE or partial IE. Simulation studies have validated that the proposed CLBC greatly outperforms two state-of-the-art modular backstepping controllers, namely HOT-ABC and MRE-HOT, in both parameter estimation and control. Further work would focus on nonlinear parameterization problems and robot control based on the proposed method.
@@ Appendix A The proof of Theorem 1
[…]
@@
$V_{\theta}=\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}$. (32)
@@
$\dot{V}_{\theta}\leq-2\kappa\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e})\tilde{\bm{\theta}}-2\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$.
@@
$\tilde{\bm{\theta}}^{T}(t)\Gamma^{-1}\tilde{\bm{\theta}}(t)\leq\tilde{\bm{\theta}}^{T}(0)\Gamma^{-1}\tilde{\bm{\theta}}(0).$
After some algebraic operations, the above result leads to
$\|\tilde{\bm{\theta}}(t)\|\leq\sqrt{\lambda_{\max}(\Gamma)/\lambda_{\min}(\Gamma)}\|\tilde{\bm{\theta}}(0)\|$
@@
$\dot{\tilde{\bm{\theta}}}_{\zeta}=-\Gamma_{\zeta}\left(\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\right),t\geq T_{\rm a}$ (34)
with $Q_{\zeta}(t,t_{\rm e})$ :=:= $H(s)[\Psi_{\zeta}(t_{\rm e})]$, in which $\Gamma_{\zeta}\in\mathbb{R}^{N_{\zeta}\times N_{\zeta}}$ is a positive-definite diagonal matrix of learning rates with respect to active channels. Let a Lyapunov function candidate be
$V_{\theta,\zeta}=\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}.$ (35)
Differentiating $V_{\theta,\zeta}$ with respect to $t$ and applying (34) to the resulting expression, one obtains
${\dot{V}}_{\theta,\zeta}\leq$ $-2\tilde{\bm{\theta}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}-2\kappa\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}$
$\leq$ $-2\kappa\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}.$
@@
$\dot{V}_{\theta,\zeta}\leq-2\kappa\sigma^{*}\tilde{\bm{\theta}}_{\zeta}^{T}{\tilde{\bm{\theta}}}_{\zeta}\leq-2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta}.$
@@
with $k_{\zeta}$ :=:= $2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\in\mathbb{R}^{+}$. Applying [88, Lemma A.3.2] to solve the above inequality, one obtains
@@
$\|\tilde{\bm{\theta}}_{\zeta}(t)\|^{2}\leq$ $\lambda_{\max}(\Gamma_{\zeta})\tilde{\bm{\theta}}_{\zeta}^{T}(t)\Gamma_{\zeta}^{-1}\tilde{\bm{\theta}}_{\zeta}(t)=\lambda_{\max}(\Gamma_{\zeta})V_{\theta,\zeta}(t)$
$\leq$ $\lambda_{\max}(\Gamma_{\zeta})V_{\theta,\zeta}(T_{\rm a})e^{-k_{\zeta}(t-T_{\rm a})},\forall t\geq T_{\rm a}$
@@
$\|\tilde{\bm{\theta}}(t)\|^{2}\leq$ $\lambda_{\max}(\Gamma)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}=\lambda_{\max}(\Gamma)V_{\theta}(t)$
$\leq$ $\lambda_{\max}(\Gamma)V_{\theta}(T_{\rm e})e^{-k_{\theta}(t-T_{\rm e})},\forall t\geq T_{\rm e}$
where $k_{\theta}$ :=:= $2\kappa\sigma^{*}\lambda_{\min}(\Gamma)$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies

### v4

@@ V-C Robustness Results
Theorem 3: For the system (4) with the additive disturbance $\bm{d}$ and Assumptions 1–2 driven by the CLBC law (6) and (23) with $\bm{x}(0)\in\Omega_{{\rm c}_{0}}$ and $\hat{\bm{\theta}}(0)\in\Omega_{{\rm c}_{\theta}}$, there exist suitable control parameters $k_{{\rm c}1}$ to $k_{{\rm c}n}$ in (5) and filtered parameters $\alpha_{1}$ to $\alpha_{n-1}$ in (16), such that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) has:
@@
    Partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$ when partial IE exists for constants $T_{\rm a}$, $\sigma\in\mathbb{R}^{+}$ and the set $\mathcal{I}$ no longer changes, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa_{1}$, $\kappa_{2}$, and $\bar{d}$;
@@
    Practical exponential stability on $t$ $\in$ $T_{\rm e},\infty)$ if IE exists for constants $T_{\rm e}$, $\sigma$ $\in$ $\mathbb{R}^{+}$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa_{1}$, $\kappa_{2}$, and $\bar{d}$.
@@ VI-A Stability and Convergence Comparisons
Set the control parameters $k_{{\rm c}1}=k_{{\rm c}2}=k_{{\rm c}3}=$ 1, $\kappa_{1}=\kappa_{2}$ == $\tau_{\rm d}=3$, $\hat{\bm{\theta}}(0)=\bm{0}$, and $\sigma=10^{-4}$, and the stable filter $H(s)=25/(s^{2}+10s+25)$ in (16). Gaussian white noise with mean 0 and standard deviation 0.001 is added to the measurements of the system states $x_{i}$. The MRE-HOT in [32] and the composite learning dynamic surface control (CL-DSC) in [17] are selected as baseline controllers, with the damping parameters $k_{\rm{d}\it i}=$ 0.1 for the MRE-HOT, the stable filter $L(s)$ == $20/(s+20)$ for the CL-DSC, and the other shared parameters being the same values as the proposed CLBC for fair comparisons.
@@
Case 1: Tracking with PE. Consider a tracking problem under PE with the reference trajectory $y_{\rm r}=1.5\sin(0.5t)$, the desired parameter $\bm{\theta}=[0.1,0.5,1.5]^{T}$, and the initial state $\bm{x}(0)$ == $0.6,0,0]^{T}$. Performance comparisons of the three controllers are depicted in Fig. 3. It is observed that the estimation error $\tilde{\bm{\theta}}$ by the CL-DSC has a steady-state error since the time derivative of each virtual control input is estimated by the filter $L(s)$ [see Fig. 3(a)], $\tilde{\bm{\theta}}$ by the MRE-HOT converges to $\bm{0}$ after running 30 s [see Fig. 3(a)] as PE is fulfilled in this case, and $\tilde{\bm{\theta}}$ by the proposed CLBC exhibits the rapid convergence to $\bm{0}$ [see Fig. 3(a)], which validates its strong learning capability. Besides, the proposed CLBC exhibits much better tracking performance than the CL-DSC and MRE-HOT [see Fig. 3(b)] because the exciting strength $\sigma_{\rm c}$ of the CLBC is monotonic non-decreasing and keeps a high level throughout [see Fig. 3(d)]. Moreover, the control inputs $u$ by the proposed CLBC and the MRE-HOT are not affected by the measurement noise throughout and are comparable after their estimation errors $\tilde{\bm{\theta}}$ converge to $\bm{0}$ [see Fig. 3(c)], but $u$ by the CL-DSC is seriously polluted by the measurement noise [see Fig. 3(c)], which results in much worse tracking performance [see Fig. 3(b)].
@@ VI-B Transient Performance Comparisons
with $u\in\mathbb{R}$, $\varphi_{1}(x_{1})=x_{1}^{2}$, $\theta=2$, and $\bm{x}(0)=[0.6,0]^{T}$. Note that for existing modular backstepping control methods, if the damping coefficients $k_{\rm{d}\it i}$ are close to 0, the transient performance of the above system can be deteriorated by the modeling error term $\Phi^{T}\tilde{\theta}$ regardless of the convergence of the estimation error $\tilde{\theta}$ [14]. The reference trajectories $y_{\rm r}$, $\dot{y}_{\rm r}$, and $\ddot{y}_{\rm r}$ are generated by $y_{\rm r}(t)$ == $\frac{a_{0}}{a(s)}[r(t)]$ with $a_{0}=1$, $a(s)=s^{3}+3s^{2}+3s+1$, and $r(t)=\sin(2t)$. Set the control parameters $k_{{\rm c}1}=k_{{\rm c}2}=$ $\kappa_{1}$ == $\kappa_{2}=$ $\tau_{\rm d}=1$, $\hat{\theta}(0)=0$, and $\sigma=10^{-4}$, and the stable filter $H(s)$ == $25/(s+25)$. The HOT adaptive backstepping control (HOT-ABC) in [31], the MRE-HOT in [32], and the proposed CLBC without the prediction error $\bm{\epsilon}$ (i.e., set the weight factor $\kappa_{1}=0$ in (23)), are selected as baseline controllers, where their shared parameters are set to be the same for fair comparisons.
@@
Performance comparisons of transient tracking and parameter convergence of the four controllers are exhibited in Fig. 5, where both the damping parameters $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ increase from $0.01$ to $0.19$ with a step size of $0.03$, and the arrows indicate the increasing direction $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$. It is observed that the transient results of the HOT-ABC and MRE-HOT are deteriorated by decreasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ [see green and red dash lines in Fig. 5(a)]. For the MRE-HOT, increasing $k_{\rm{d}\it 1}$ and $k_{\rm{d}\it 2}$ reduces the initial oscillations of the tracking error $|e_{1}|$ [see green dash lines in Fig. 5(a)] but also slows down parameter convergence [see green dash lines in Fig. 5(b)]. For the CLBC without $\bm{\epsilon}$, the tracking error $|e_{1}|$ diverges after $t=0.7$ s due to the absence of $\bm{\epsilon}$ [see purple dash line in Fig. 5(a)]. The proposed CLBC provides the convergence of $|e_{1}|$ to 0 with the smallest initial oscillations [see Fig. 5(a)] and achieves the fastest parameter convergence [see Fig. 5(b)], which implies that: 1) the transient tracking and parameter convergence of the HOT-ABC and MRE-HOT are sensitive to damping parameters; 2) the transient performance of the proposed CLBC is enhanced by introducing the extra prediction error $\bm{\epsilon}$ in (22) without resorting to nonlinear damping terms; 3) The combination of two prediction errors $\bm{\xi}$ in (19) and $\bm{\epsilon}$ in (22) accelerates parameter convergence.
@@ VII Conclusions
This paper has presented a feasible modular backstepping control strategy named CLBC for strict-feedback uncertain nonlinear systems. The proposed composite learning HOT allows the exact implementation of the high-order time derivatives of parameter estimates and the offset of modeling errors, such that the transient performance can be guaranteed without resorting to nonlinear damping terms or high control gains. The proposed algorithm of staged exciting strength maximization ensures that the exciting strength is monotonically non-decreasing at each excitation stage, thereby enabling exponential stability of the closed-loop system with parameter convergence under the much weaker condition of IE or partial IE. Simulation studies have validated that the proposed CLBC greatly outperforms two state-of-the-art modular backstepping controllers, namely HOT-ABC and MRE-HOT, in both parameter estimation and control. Further work would focus on nonlinear parameterization problems and robot control based on the proposed method.
@@ Appendix A The proof of Theorem 1
To facilitate the proof, set the weight factor $\kappa_{1}=1$ in (23), and this will also be used in the proofs of Theorems 2–3.
@@
$\dot{\tilde{\bm{\theta}}}=-\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}-\kappa_{2}Q(t,t_{\rm e})\tilde{\bm{\theta}}$. (31)
@@
$V_{\theta}=\tilde{\bm{\theta}}^{T}\tilde{\bm{\theta}}$. (32)
@@
$\dot{V}_{\theta}\leq-2\kappa_{2}\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e})\tilde{\bm{\theta}}-2\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$.
@@
$\|\tilde{\bm{\theta}}(t)\|\leq\|\tilde{\bm{\theta}}(0)\|$
@@
$\dot{\tilde{\bm{\theta}}}_{\zeta}=-\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}-\kappa_{2}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta},t\geq T_{\rm a}$ (34)
with $Q_{\zeta}(t,t_{\rm e})$ :=:= $H(s)[\Psi_{\zeta}(t_{\rm e})]$. Choose
$V_{\theta,\zeta}=\tilde{\bm{\theta}}_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}$ (35)
as a Lyapunov function candidate. Differentiating $V_{\theta,\zeta}$ with respect to $t$ and applying (34), one obtains
${\dot{V}}_{\theta,\zeta}\leq$ $-2\tilde{\bm{\theta}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}-2\kappa_{2}\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}$
$\leq$ $-2\kappa_{2}\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}.$
@@
$\dot{V}_{\theta,\zeta}\leq-2\kappa_{2}\sigma^{*}\tilde{\bm{\theta}}_{\zeta}^{T}{\tilde{\bm{\theta}}}_{\zeta}.$
@@
with $k_{\zeta}$ :=:= $2\kappa_{2}\sigma^{*}\in\mathbb{R}^{+}$. Applying [42, Lemma A.3.2] to solve the above inequality, one obtains
@@
$\|\tilde{\bm{\theta}}_{\zeta}(t)\|^{2}=V_{\theta,\zeta}(t)\leq V_{\theta,\zeta}(T_{\rm a})e^{-k_{\zeta}(t-T_{\rm a})},\forall t\geq T_{\rm a}$
@@
$\|\tilde{\bm{\theta}}(t)\|^{2}=V_{\theta}(t)\leq V_{\theta}(T_{\rm e})e^{-k_{\theta}(t-T_{\rm e})},\forall t\geq T_{\rm e}$
where $k_{\theta}$ :=:= $2\kappa_{2}\sigma^{*}$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies